OpenStudy (anonymous):

Help with #6 pls!!! It's substitution with system of equations! I will fan and medal!!

OpenStudy (anonymous):

First divide 8x-8y=-24 by 8 to get x-y=-3 Then add that to 3x+y=23 to isolate x 3x+y=23 X-y=-3 4x=20 x=5 Then plug in x to find the y value

OpenStudy (anonymous):

Thank u so much!!

OpenStudy (anonymous):

Welcome! :)

OpenStudy (anonymous):

Which equation would I plug x in though?

OpenStudy (anonymous):

@angeltears

OpenStudy (anonymous):

Either one is fine

OpenStudy (anonymous):

Would you like me to show you how?

OpenStudy (anonymous):

So let's try to plug it into the second one 3x+y=23 3(5)+y=23 15+y=23 Subtract 15 on both sides Y=8

OpenStudy (anonymous):

8x-8y=-24 8(5)-8y=-24 40-8y=-24 Subtract 40 on both sides -8y=-64 Divide by -8 Y=8

OpenStudy (anonymous):

Ok thx u